# Adaptive Gradient Calibration for Single-Positive Multi-Label Learning in Remote Sensing Image Scene Classification

### v1

@@ III-C Early Learning Detection
Pseudo-labelling methods typically require an initial warm-up phase, whose duration, often determined empirically, critically impacts performance. In SPML, premature pseudo-label initiation introduces additive noise, whereas excessive delay risks overfitting to inherent subtractive label noise, thus limiting false negative detection. Optimal pseudo-labelling activation is crucial for model performance. Monitoring performance metrics on a clean validation set is ideal, as model performance typically improves, plateaus, then declines (indicating overfitting) [40, 41]. The plateau, detectable by the slope of performance metrics, identifies the optimal time to activate pseudo-labelling [22]. However, SPML generally lacks clean validation data, typically offering only single positive labels, akin to the training set. While standard metrics on noisy sets might differ, we demonstrate that the mean average precision (mAP) computed on the noisy validation set mirrors the behavior of mAP* computed on the clean validation set, showing approximately consistent trends throughout training.
The mAP is the macro average of average precision (AP) for each class:
$\text{mAP}=\frac{1}{C}\sum_{c=1}^{C}{\text{AP}_{c}},$ (11)
where $\text{AP}_{c}=\int_{0}^{1}{\text{Prec}_{c}\left(\text{Rec}_{c}\right)d\text{Rec}_{c}}$. For a clean validation set, recall $\text{Rec}^{*}_{c}$ and precision $\text{Prec}^{*}_{c}$ are defined as follows:
$\begin{split}\text{Rec}^{*}_{c}=\frac{\text{TP}_{c}}{\text{P}_{c}}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FN}_{c}},\\ \text{Prec}^{*}_{c}=\frac{\text{TP}_{c}}{\text{PP}_{c}}=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FP}_{c}},\end{split}$ (12)
where $\text{TP}_{c}$, $\text{FP}_{c}$, and $\text{FN}_{c}$ are the true positives, false positives, and false negatives for class $c$, respectively, while $\text{P}_{c}=\text{TP}_{c}+\text{FN}_{c}$ and $\text{PP}_{c}=\text{TP}_{c}+\text{FP}_{c}$ are the real and predicted positives. In SPML, flipped positives are $\text{F}_{c}=\beta_{c}\cdot\text{P}_{c}$, where $\beta_{c}\in(0,1)$ is the proportion of flipped positives. Let $\text{PF}_{c}=\alpha_{c}\cdot\text{F}_{c}$ be the proportion of correctly classified flipped positives, where $\alpha_{c}\in[0,1]$. Thus, noisy validation metrics $\text{Rec}_{c}$ and $\text{Prec}_{c}$ are functions of their clean counterparts:
$\begin{split}\text{Rec}_{c}&=\frac{\text{TP}_{c}-\text{PF}_{c}}{\text{P}_{c}-\text{F}_{c}}=\frac{\text{Rec}^{*}_{c}-\alpha_{c}\cdot\beta_{c}}{1-\beta_{c}},\\ \text{Prec}_{c}&=\frac{\text{TP}_{c}-\text{PF}_{c}}{\text{PP}_{c}}=\text{Prec}^{*}_{c}\cdot\left(1-\frac{\alpha_{c}\cdot\beta_{c}}{\text{Rec}^{*}_{c}}\right).\end{split}$ (13)
These metrics depend on the label noise flip rate $\beta_{c}$ and the classifier’s behavior on flipped labels $\alpha_{c}$, which can vary, altering the noisy validation dynamic from the clean one. However, assumptions about $\alpha_{c}$ based on noise type allow us to draw some conclusions. We consider two SPML noise types: Random and Dominant. In both cases, we will demonstrate below that the expected noisy $\text{AP}_{c}$ for class $c$ relates to its true counterpart $\text{AP}_{c}^{*}$ via a class-specific scaling factor $\omega_{c}$:
$\mathbb{E}[\text{AP}_{c}]=\omega_{c}\cdot\text{AP}_{c}^{*}.$ (14)
Consequently, the expected noisy mAP, can be expressed relative to the true mAP* as follows:
$\mathbb{E}[\text{mAP}]=\frac{1}{C}\sum_{c=1}^{C}{\omega_{c}\cdot\text{AP}_{c}^{*}}=\bar{\omega}\cdot\text{mAP}^{*}+\text{Cov}\left(\omega,\text{AP}^{*}\right),$ (15)
where $\bar{\omega}=\frac{1}{C}\sum_{c=1}^{C}{\omega_{c}}$ is the mean scaling factor across classes, and $\text{Cov}\left(\omega,\text{AP}^{*}\right)$ denotes the covariance between the class-specific scaling factors and the true APs. The sign and magnitude of this covariance term determine the deviation from a simple mean-scaling relationship.
III-C1 Random SPML
In random SPML, single positive labels are chosen uniformly at random from the set of true positives for each instance. This type of noise is generally considered instance-independent [42]. While the trained model may inevitably fit this noise in the training labels, it cannot meaningfully generalize to it or predict the noisy labels in the validation set. Under this assumption, the proportion of correctly classified flipped positives, $\alpha_{c}$, on average equals the recall for that class, i.e., $\mathbb{E}\left[\alpha_{c}\right]\approx\text{Rec}^{*}_{c}$, from which we can derive the following noisy validation metrics:
$\mathbb{E}\left[\text{Rec}_{c}\right]\approx\text{Rec}^{*}_{c},\quad\mathbb{E}\left[\text{Prec}_{c}\right]\approx\left(1-\beta_{c}\right)\cdot\text{Prec}^{*}.$ (16)
Substituting these into the AP definition, we find that the expected noisy AP for each class is scaled down by the non-flip rate:
$\mathbb{E}[\text{AP}_{c}]\approx(1-\beta_{c})\cdot\text{AP}_{c}^{*}.$ (17)
Here, the scaling factor is $\omega_{c}=(1-\beta_{c})$. Substituting this into the general mAP relationship in (15), we get the following:
$\begin{split}\mathbb{E}[\text{mAP}]&\approx\frac{1}{C}\sum_{c=1}^{C}{(1-\beta_{c}})\cdot\text{AP}_{c}^{*}<\text{mAP}^{*},\\ \mathbb{E}[\text{mAP}]&\approx(1-\bar{\beta})\cdot\text{mAP}^{*}-\text{Cov}\left(\beta,\text{AP}^{*}\right),\end{split}$ (18)
where $\bar{\beta}$ is the mean flip rate across classes. This implies that the expected noisy mAP will be lower than the true mAP*. The covariance term $\text{Cov}\left(\beta,\text{AP}^{*}\right)$ captures how the class-specific flip rates correlate with their true performances. In general, we can expect that better-performing classes experience less flipping, implying negative covariance and therefore a mitigation effect on the mean-scaling relationship in (18), i.e., $(1-\bar{\beta})\cdot\text{mAP}^{*}\leq\mathbb{E}[\text{mAP}]<\text{mAP}^{*}$.
III-C2 Dominant SPML
In contrast to random SPML, dominant SPML belongs to the category of instance-dependent label noise [42]. Here, the single-positive label is consistently the “dominant” one among the true positives. This allows the model to potentially learn this label-noise model, which can be reflected in the noisy validation mAP later in training. This results in three consecutive trends during learning [22]: i) an early learning phase where $\mathbb{E}\left[\alpha_{c}\right]\approx\text{Rec}^{*}_{c}$ still holds, yielding similar results to random SPML; ii) a transition phase; and iii) a label-noise model fitting phase. Once the model learns the label-noise behavior, the expected proportion of correctly classified flipped positives approaches zero, as the model learns to ignore non-dominant true positives, i.e., $\mathbb{E}\left[\alpha_{c}\right]\to 0$. This implies the following expected noisy validation metrics:
$\mathbb{E}[\text{Rec}_{c}]\approx\frac{\text{Rec}_{c}^{*}}{1-\beta_{c}},\quad\mathbb{E}[\text{Prec}_{c}]\approx\text{Prec}_{c}^{*},$ (19)
which leads to the following expected AP relationship:
$\mathbb{E}[\text{AP}_{c}]\approx\frac{1}{1-\beta_{c}}\cdot\text{AP}_{c}^{*}.$ (20)
In this case, the scaling factor is $\omega_{c}={1}/(1-\beta_{c})$. The corresponding mAP relationship is as follows:
$\small\begin{split}\text{mAP}^{*}&<\mathbb{E}[\text{mAP}]=\frac{1}{C}\sum_{c=1}^{C}{\frac{1}{1-\beta_{c}}\cdot\text{AP}_{c}^{*}}<1,\\ \mathbb{E}[\text{mAP}]&\approx\left(\frac{1}{C}\sum_{c=1}^{C}{\frac{1}{1-\beta_{c}}}\right)\text{mAP}^{*}+\text{Cov}\left(\frac{1}{1-\beta},\text{AP}^{*}\right).\end{split}$ (21)
This indicates that the expected noisy mAP will be higher than the true $\text{mAP}_{c}^{*}$. The covariance term captures if classes with higher true APs are affected by the scaling. In general, we expect highly dominant classes to perform better than others. This leads to a negative covariance value, i.e., $\text{mAP}^{*}<\mathbb{E}[\text{mAP}]\leq\left(\frac{1}{C}\sum_{c=1}^{C}{\frac{1}{1-\beta_{c}}}\right)\text{mAP}^{*}$. Consequently, noisy $\text{AP}_{c}$ are initially lower than $\text{AP}_{c}^{*}$ but later they increase their value. The noisy $\text{AP}_{c}$ approximately remain proportional to $\text{AP}_{c}^{*}$ during these phases, with the “transition phase” marking a departure from this simple scaling, possibly shifting the plateau’s appearance between them.
III-C3 Robust Plateau Detection using Teacher Model
@@
Figure 3: True and noisy validation mAP trends of the teacher model during the training with AN on the refined BigEarthNet [14] dataset: (a) random SPML, and (b) dominant SPML.
As demonstrated in (18) and (21), although a proportional relationship generally holds between the noisy mAP and the true $\text{mAP}^{*}$, the bias term (i.e., the covariance) can introduce deviations from this relationship. However, we argue that this bias term rarely compromises the analysis, especially in the early learning stage. At the beginning, the model primarily fits correctly labelled samples [21], leading to a weak correlation between $\beta_{c}$ and $\text{AP}_{c}^{*}$. As a result, the scaling relationship predominantly governs (18) and (21). Later, when the $\text{AP}_{c}^{*}$ scores plateau, the covariance term’s contribution to the slope of the noisy mAP becomes negligible, as it exhibits little to no variation. Together, these factors enable the noisy mAP to closely mirror the true performance plateau, making it a reliable indicator for identifying the end of the early learning phase.
We propose monitoring the teacher model’s epoch-wise noisy validation mAP score to ensure a stable and reliable detection. The teacher model is maintained as an EMA of the student model’s weights. It acts as a low-pass filter, smoothing the student’s unstable learning dynamics with fluctuations caused by noisy gradients or erratic updates. As shown in Fig. 3, the noisy and true mAP scores computed from the teacher model closely follow similar plateau trends, consistent with the relationships established in (18) and (21). This smoothed behavior allows for a more accurate and robust detection of the optimal epoch for pseudo-labeling.
In practice, we online select the epoch that achieves the highest noisy validation mAP score from the teacher model, provided there is no significant improvement within the following $b_{e}$ epochs. Once this optimal epoch is detected, the model weights are automatically restored from that epoch, and pseudo-labeling is initiated concurrently with GC.
@@ III-D Dual-EMA Pseudo-Label Generation
As discussed in Section III-B, GC relies on high-quality pseudo-labels. Since it is triggered after early learning detection on the teacher model, the teacher’s predictions are a natural choice. However, due to slow weight updates, the teacher model may exhibit a delay in incorporating recently acquired knowledge after GC begins. Therefore, we propose to combine predictions from both teacher and student models. Specifically, we apply EMA to the student’s predictions, using a smaller smoothing factor than that for the teacher model, to avoid overfitting to its own outputs while maintaining adaptability.
@@
with $\theta^{T}(e)$ and $\theta^{S}(e)$ being the model weights of the teacher and student models at epoch $e$, respectively. We set the smoothing coefficient $\beta_{t}=0.999$ following [43]. Then, we compute the temporarily EMA-smoothed student predictions at the $e$th iteration with the coefficient $\beta_{s}\in[0,1)$ as follows:
@@
We empirically set $\beta_{s}=0.8$ to guarantee the update speed of the student model’s prediction. Then, the pseudo label $\mathbf{t}_{i}(e)$ used for gradient calibration is computed as the combination of the teacher and student predictions as follows:
@@ III-E Mixup
To further enhance robustness against label noise and improve generalization, we incorporate Mixup [15] into the GC stage. For each batch, we randomly sample a pairing input $(\mathbf{x}_{j},\mathbf{y}_{j},\mathbf{t}_{j})$ for every original input $(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{t}_{i})$, and compute:
@@
IV Experiments
This section details the experimental setup employed to evaluate the proposed AdaGC approach against a comprehensive set of state-of-the-art (SOTA) SPML methods. We first describe the datasets utilized and the related specific SPML noise types considered. Subsequently, we outline the experimental setup, including hyperparameter setting, evaluation metrics, model architectures, and comparison methods considered in this study. Finally, we present the comparison results and an in-depth ablation study of AdaGC.
IV-A Datasets
We evaluate the proposed method on two multi-label RS benchmark datasets characterized by different modalities and spatial resolutions: refined BigEarthNet (reBEN) [44, 45, 14], derived from SAR and multispectral Sentinel-1/2 imagery with a spatial resolution of 10m, and AID-multilabel [13], collected from Google Earth aerial RGB imagery with resolutions ranging from 0.5m to 8m per pixel. reBEN focuses on land cover classification, while AID-multilabel includes both land cover categories and object-level semantic information. Due to the availability of CORINE Land Cover (CLC) [46] pixel-wise masks for reBEN, we were able to simulate the two types of label noise described in Section III-C. In contrast, AID’s high spatial resolution introduces greater semantic complexity, making it challenging to determine a dominant class. Therefore, we simulate only Random single-positive label noise on AID.
IV-A1 Refined BigEarthNet
It contains 549,488 Sentinel-1/2 image pairs of size of $120\times 120$ pixels covering 10 European countries, each aligned with a pixel-level CLC reference map. Multi-label annotations are derived from these maps, resulting in 19 land cover classes (see Table I). Following the splits in [14], we use a 2:1:1 ratio for training, validation, and testing. Each input has 14 channels by concatenating 2 Sentinel-1 and 12 Sentinel-2 L2A bands. Leveraging the reference maps, on this dataset we simulate both types of SPML label noise (Random and Dominant). Dataset statistics for the two cases are provided in Table I.
@@ IV-A1 Refined BigEarthNet
Random and Dominant SPML Label Noise. Macro
and Micro Averages Are Provided, as Well as
the Average Number of True Positive
Labels per Image
@@
Urban fabric 0.73 0.84 33626
Industrial or commercial units 0.72 0.91 6398
Arable land 0.63 0.51 88236
Permanent crops 0.70 0.74 15190
Pastures 0.66 0.62 42968
Complex cultivation patterns 0.73 0.77 51034
Land principally occupied by agriculture, with significant areas of natural vegetation 0.73 0.86 60127
Agro-forestry areas 0.64 0.54 15082
Broad-leaved forest 0.69 0.65 64164
Coniferous forest 0.66 0.54 76344
Mixed forest 0.70 0.67 78791
Natural grassland and sparsely vegetated areas 0.69 0.82 6984
Moors, heathland and sclerophyllous vegetation 0.70 0.70 6329
Transitional woodland, shrub 0.70 0.85 64198
@@
Inland waters 0.67 0.73 30466
Marine waters 0.08 0.05 35914
@@
Average #labels/image 2.9

### v2

@@ III-C Early Learning Detection
Pseudo-labelling methods typically require an initial warm-up phase to mitigate the trade-off between additive noise (premature initiation) and subtractive noise overfitting (excessive delay). Ideally, the optimal activation time is identified by the performance plateau on a clean validation set [Arpit_2017_ICML, Zhang_2021_CommACM, liu_aio2_2024]. In SPML, where clean validation data is unavailable, we propose monitoring the mean average precision (mAP) on the noisy validation set. Below, we provide a formal justification for this criterion by analyzing the bounds and stationarity of the noisy metric.
III-C1 Theoretical Framework
@@
Figure 3: True and noisy validation mAP trends of the teacher model during training on refined BigEarthNet [clasen_reben_2025_arxivonly]: (a) random SPML, (b) dominant SPML. The trends align with Propositions 1 and 2, and confirm the validity of the stationarity theorem.
Let $\text{Rec}^{*}_{c}$ and $\text{Prec}^{*}_{c}$ denote the clean recall and precision for class $c$. In SPML, the set of real positives $\text{P}_{c}$ is reduced by the flipped positives $\text{F}_{c}=\beta_{c}\text{P}_{c}$, where $\beta_{c}\in(0,1)$ is the noise rate. Defining $\text{PF}_{c}=\alpha_{c}\cdot\text{F}_{c}$, where $\alpha_{c}\in[0,1]$, as the proportion of correctly classified flipped positives, the noisy metrics relate to the clean ones as follows:
$\begin{split}\text{Rec}_{c}&=\frac{\text{TP}_{c}-\text{PF}_{c}}{\text{P}_{c}-\text{F}_{c}}=\frac{\text{Rec}^{*}_{c}-\alpha_{c}\beta_{c}}{1-\beta_{c}},\\ \text{Prec}_{c}&=\frac{\text{TP}_{c}-\text{PF}_{c}}{\text{PP}_{c}}=\text{Prec}^{*}_{c}\left(1-\frac{\alpha_{c}\beta_{c}}{\text{Rec}^{*}_{c}}\right),\end{split}$ (11)
where $\text{TP}_{c}$ and $\text{PP}_{c}$ are the true and predicted positives, respectively. The mAP is the macro average of average precision (AP) for each class:
$\text{mAP}=\frac{1}{C}\sum_{c=1}^{C}{\text{AP}_{c}},$ (12)
where $\text{AP}_{c}=\int_{0}^{1}{\text{Prec}_{c}\left(\text{Rec}_{c}\right)d\text{Rec}_{c}}$.
Proposition 1 (Random SPML Lower Bound). Under instance-independent noise, the expected noisy mAP is a lower bound of the clean $\text{mAP}^{*}$.
Proof: In Random SPML, the noise is independent of the input features [Perantoni_2022_TGRS]. Thus, the model cannot generalize to the random flips, implying $\mathbb{E}[\alpha_{c}]\approx\text{Rec}^{*}_{c}$. Substituting this into (11) yields $\mathbb{E}[\text{AP}_{c}]\approx(1-\beta_{c})\text{AP}_{c}^{*}$. The expected noisy mAP becomes:
$\mathbb{E}[\text{mAP}]\approx(1-\bar{\beta})\text{mAP}^{*}-\text{Cov}(\beta,\text{AP}^{*}),$ (13)
where $\bar{\beta}$ is the mean flip rate. Since better-performing classes generally experience less flipping (negative covariance), $\mathbb{E}[\text{mAP}]<\text{mAP}^{*}$. $\square$
Fig. 3(a) shows the trends observed while training on the refined BigEarthNet [clasen_reben_2025_arxivonly] dataset with random single-positive labels, which are aligned with our theoretical analysis.
Proposition 2 (Dominant SPML Upper Bound). Under instance-dependent dominant noise, the expected noisy mAP is an upper bound of the clean $\text{mAP}^{*}$.
Proof: In Dominant SPML, the model learns the label-noise model [Perantoni_2022_TGRS], ignoring non-dominant true positives ($\mathbb{E}[\alpha_{c}]\to 0$). Eq. (11) simplifies to $\mathbb{E}[\text{Rec}_{c}]\approx\frac{\text{Rec}_{c}^{*}}{1-\beta_{c}}$ and $\mathbb{E}[\text{Prec}_{c}]\approx\text{Prec}_{c}^{*}$, yielding $\mathbb{E}[\text{AP}_{c}]\approx\frac{1}{1-\beta_{c}}\text{AP}_{c}^{*}$. The noisy mAP becomes:
$\mathbb{E}[\text{mAP}]\approx\overline{\left(\frac{1}{1-\beta}\right)}\text{mAP}^{*}+\text{Cov}\left(\frac{1}{1-\beta},\text{AP}^{*}\right).$ (14)
Assuming dominant classes perform better (negative covariance), $\text{mAP}^{*}<\mathbb{E}[\text{mAP}]$. $\square$
Generally, Dominant SPML results in three consecutive trends during learning [liu_aio2_2024]: i) an early learning phase where $\mathbb{E}\left[\alpha_{c}\right]\approx\text{Rec}^{*}_{c}$ still holds, yielding similar results to random SPML; ii) a transition phase; and iii) a label-noise model fitting phase. Fig. 3(b) plots the trends observed while training on the refined BigEarthNet [clasen_reben_2025_arxivonly] dataset with dominant single-positive labels, showing this three-phase phenomenon.
III-C2 Adaptive Triggering Mechanism
We now address the validity of using the noisy metric for early stopping.
Theorem 1 (Stationarity Preservation). The stationary points of the expected noisy mAP coincide with those of the clean $\text{mAP}^{*}$ in the late early-learning phase.
Proof: Let $t$ denote the training epoch. Combining Propositions 1 and 2, the noisy metric is a linear transformation of the clean metric: $\mathbb{E}[\text{mAP}(t)]=\bar{\omega}\text{mAP}^{*}(t)+\text{Cov}(\omega,\text{AP}^{*}(t))$. Taking the temporal derivative:
$\frac{\partial\mathbb{E}[\text{mAP}]}{\partial t}=\bar{\omega}\frac{\partial\text{mAP}^{*}}{\partial t}+\frac{\partial}{\partial t}\text{Cov}(\omega,\text{AP}^{*}(t)).$ (15)
The noise parameters $\omega$ are constant. Near the performance plateau ($t\to t^{*}$), class-wise learning dynamics tend to stabilize, making the covariance term approximately constant at that time ($\frac{\partial}{\partial t}\text{Cov}\approx 0$) Consequently:
$\frac{\partial\mathbb{E}[\text{mAP}]}{\partial t}\propto\frac{\partial\text{mAP}^{*}}{\partial t}.$ (16)
Thus, identifying the peak of the noisy mAP is mathematically equivalent to identifying the peak of the clean mAP. $\square$
Implementation: To mitigate stochastic fluctuations in the single-run estimates of $\mathbb{E}[\text{mAP}]$, we monitor the mAP of the teacher model (maintained via EMA), which acts as a low-pass filter on the model’s weights (see Fig. 6). We select the epoch maximizing the teacher’s noisy validation mAP (with a patience of $b_{e}$ epochs) to initiate pseudo-labeling.
@@ III-D Dual-EMA Pseudo-Label Generation
As discussed in Section III-B, GC relies on high-quality pseudo-labels. The teacher’s predictions are a natural choice, since the early learning detection is on the teacher model. However, due to slow weight updates, the teacher model may exhibit a delay in incorporating recently acquired knowledge after GC begins. Therefore, we propose to combine predictions from both teacher and student models. Specifically, we apply EMA to the student’s predictions with a smaller smoothing factor than that for the teacher model, to avoid overfitting to its own outputs while maintaining adaptability.
@@
with $\theta^{T}(e)$ and $\theta^{S}(e)$ being the model weights of the teacher and student models at iteration $e$. We set the smoothing coefficient $\beta_{t}=0.999$ following [tarvainen_mean_2017]. Then, the temporarily EMA-smoothed student predictions are as follows:
@@
We empirically set $\beta_{s}=0.8$ to guarantee the update speed of the student’s predictions. Then, the pseudo-label $\mathbf{t}_{i}(e)$ used for gradient calibration is computed as the combination of the teacher and student predictions as follows:
@@ III-E Mixup
To further enhance robustness against label noise and improve generalization, we incorporate Mixup [Zhang_2018_ICLR] into the GC stage. For each batch, we randomly sample a pairing input $(\mathbf{x}_{j},\mathbf{y}_{j},\mathbf{t}_{j})$ for every original input $(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{t}_{i})$, and compute:
@@
III-F Complexity Analysis
To assess practical feasibility, we analyze the computational complexity of AdaGC relative to a standard classification baseline. First, one can note that AdaGC is a training-only framework, i.e., the dual-EMA module is discarded after training. Consequently, the inference complexity (parameters and FLOPs) is identical to the student backbone, ensuring no additional latency or memory overhead during deployment.
Regarding training efficiency, a standard iteration incurs a cost of approximately $3\times$ the forward pass FLOPs ($1\text{F}$ forward and $\sim 2\text{F}$ backward). Instead, AdaGC processes two views during the GC stage: the original batch and a Mixup-augmented batch. However, the original batch is utilized exclusively for generating robust pseudo-labels via the student and teacher networks, which requires only forward passes, while the model optimization relies on the Mixup batch. Therefore, the computationally intensive backward pass is strictly necessary only for this view. This design results in a total theoretical cost of $\sim 5\text{F}$ per iteration, representing a moderate $\sim 66\%$ overhead compared to the baseline ($3\text{F}$) exclusively during the GC stage. Note that Mixup is simply a linear combination of two images. Thus, its additional FLOPs are mathematically negligible compared to the rest of the pipeline.
In terms of memory complexity, the overhead is modest. The teacher EMA maintains only a second copy of the backbone parameters (not activations), contributing less than 10% of total memory in typical large-scale models. The temporally smoothed student predictions are stored as image-wise labels, which also require minimal memory. Taken together, both FLOPs and memory considerations indicate that AdaGC is computationally feasible for large-scale remote sensing applications.
@@ III-F Complexity Analysis
Random and Dominant SPML Label Noise
@@
Urban fabric 0.73 0.84 33,626
Industrial or commercial units 0.72 0.91 6,398
Arable land 0.63 0.51 88,236
Permanent crops 0.70 0.74 15,190
Pastures 0.66 0.62 42,968
Complex cultivation patterns 0.73 0.77 51,034
Land principally occupied by agriculture, with significant areas of natural vegetation 0.73 0.86 60,127
Agro-forestry areas 0.64 0.54 15,082
Broad-leaved forest 0.69 0.65 64,164
Coniferous forest 0.66 0.54 76,344
Mixed forest 0.70 0.67 78,791
Natural grassland and sparsely vegetated areas 0.69 0.82 6,984
Moors, heathland and sclerophyllous vegetation 0.70 0.70 6,329
Transitional woodland, shrub 0.70 0.85 64,198
@@
Inland waters 0.67 0.73 30,466
Marine waters 0.08 0.05 35,914
@@
Average number of true positive labels per image 2.9
IV Experiments
This section presents the experimental setups, comparative results, and a comprehensive ablation study of AdaGC.
IV-A Datasets
We evaluate the proposed method on two multi-label RS benchmark datasets with different modalities and resolutions: refined BigEarthNet (reBEN) [Sumbul_ben_2019_IGARSS, sumbul_BigEarthNet-mm_2021, clasen_reben_2025_arxivonly], derived from SAR and multispectral Sentinel-1/2 imagery at 10m resolution, and AID-multilabel [hua_relation_2020], derived from high-resolution RGB images at 0.5–8m resolution. While reBEN focuses on land cover classification, AID-multilabel includes both land cover categories and object-level semantic information.
IV-A1 reBEN
TABLE II: AID Multilabel Classes and Related Flip Rates $\beta_{c}$ on
the Training Subset in the Two Cases of Random
and Manual SPML Label Noise
Class Name Random $\beta_{c}$ Manual $\beta_{c}$ Support
Airplane 0.83 0.80 59
Bare-soil 0.80 0.78 882
Buildings 0.82 0.81 1,316
Cars 0.85 0.85 1,210
Chaparral 0.76 0.85 79
Court 0.86 0.78 203
Dock 0.94 0.83 166
Field 0.66 0.57 127
Grass 0.78 0.80 1,368
Pavement 0.83 0.84 1,367
Sand 0.41 0.48 153
Sea 0.77 0.67 133
Ship 0.90 0.89 181
Tanks 0.82 0.73 66
Trees 0.79 0.80 1,467
Water 0.83 0.84 510
Micro average 0.81 0.81
Macro average $\bar{\beta}$ 0.79 0.77
Average number of true positive labels per image 5.2
It contains 549,488 Sentinel-1/2 image pairs of size of $120\times 120$ pixels covering 10 European countries, each aligned with a pixel-level CORINE Land Cover (CLC) reference map. Multi-label annotations are derived from these maps, resulting in 19 land cover classes (see Table I). Following the splits in [clasen_reben_2025_arxivonly], we use a 2:1:1 ratio for training, validation, and testing. Each input has 14 channels by concatenating 2 Sentinel-1 and 12 Sentinel-2 L2A bands. Leveraging the reference maps, on this dataset we simulate both types of SPML label noise described in Section III-C (Random and Dominant). The Dominant case better reflects real-world single-positive annotations, as low-resolution images contain limited details, where the prevailing class is more likely to be observed while minor classes are ignored. Dataset statistics for the two cases are provided in Table I.